OpenStudy (anonymous):

write equation in slope-intercept form 6x-4y=8

OpenStudy (anonymous):

Slope intercept form is just \(y=mx+b\). So just solve for \(y\).

OpenStudy (anonymous):

What is the m and b in the equation 6x+4y=8

OpenStudy (anonymous):

Get Y alone first: \[(-4)y=(-6)x+8\] Now like Wio said solve for y: \[((-4)y)/(-4)=(-6/-4)x+(8/-4)\] Which simplifies to: \[y=(3/2)x-2\] m=(3/2) b=(-2) I suggest reviewing a bit more before attacking these problems since its just concepts now.

OpenStudy (anonymous):

Well how come its negative if all of the numbers are positive in the equation

OpenStudy (anonymous):

Hello ?